A long solenoid has 500 turns. When a current of 2 A is passed through it, the


resulting magnetic flux linked with each turn of the solenoid is 4×10


−3 Wb. find


self-inductance.

1
Expert's answer
2022-04-28T16:41:58-0400
"\phi=NLI"

"L=\frac{\phi}{IN}=\frac{4*10^{-3}}{2*500}=4*10^{-6}\:\rm H"